OpenStudy (anonymous):

how to solve (2x-5)^2+(x+7)^2=2x+3^2 using the quadratic formula

OpenStudy (anonymous):

first get it into ax^2+bx+c=0

OpenStudy (anonymous):

(2x-5)^2+(x+7)^2-(2x+3)^2=0 (4x^2+25)+(x^2+49)+(4x^2+9)=0 9x^2+83=0 Where would the bx come from? @juliannescriv

OpenStudy (anonymous):

bx would just be 0x

OpenStudy (anonymous):

x=\[\sqrt{166i}\]